Grade 11Physical Chemistry

In order to prepare 1 litre normal solution of KMnO4, how many grams of KMnO4 are required if the solution is to be used in acid medium for oxidationOptions: 158 g31.6 g62 g790 g

akash
10 Years agoGrade 11

2 Answers

Bhavya
10 Years ago
Dear student
In order to prepare 1 normal solution(1 litre)
N = (mass/Equivalent mass)*1/Volume of solution in litre
1= mass/(158/5) * 1/1 litre { equivalent mass = molar mass/n-factor}
mass = 158/5 = 31.6 g
Rishabh
8 Years ago
Volume and normality is 1 so mass will be equal to equivalent mass that is. 158/5=31.6 so your answer is 31.6